MONT 100N – Modeling the Environment First Writing Assignment – Stage 1

The first of the larger paper assignments this semester will deal with one of two particular points from the Papal Encyclical Letter Laudato Si’. You will have a choice which of these passages you want to address.

The point of this first assignment is for you to show that you can accurately and succinctly present information from a source, without taking a side either way on whether you agree with it or not. Your work should show that you have done some careful reading– and some serious thinking–about the passage you choose. In developing your submission, you should take into account the general frameworks for constructing essays presented in They Say, I Say. (Taking a side is the “I say” part and it will come later!)

Your paper should be about 2 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx or Google Docs file to jlittle@holycross.edu by 5:00pm on Friday, September 8.

The passages – Choose one of these.

1. (From §§183 and 186.) We need to stop thinking in terms of “interventions” to save the environment in favour of policies developed and debated by all interested parties. The participation of the latter also entails being fully informed about such projects and their different risks and possibilities; this includes not just preliminary decisions but also various follow-up activities and continued monitoring. Honesty and truth are needed in scientific and political discussions; these should not be limited to the issue of whether or not a particular project is permitted by law. … The Rio Declaration of 1992 states that ‘where there are threats of serious or irreversible damage, lack of full scientific certainty shall not be used as a pretext for postponing cost-effective measures’ which prevent environmental degradation. This precautionary principle makes it possible to protect those who are most vulnerable and whose ability to defend their interests and to assemble incontrovertible evidence is limited. If objective information suggests that serious and irreversible damage may result, a project should be halted or modified, even in the absence of indisputable proof. Here the burden of proof is effectively reversed, since in such cases objective and conclusive demonstrations will have to be brought forward to demonstrate that the proposed activity will not cause serious harm to the environment or to those who inhabit it.

2. (§116. Note: This one is more challenging.) Modernity has been marked by an excessive anthropocentrism which today, under another guise, continues to stand in the way of shared understanding and of any effort to strengthen social bonds. The time has come to pay renewed attention to reality and the limits it imposes; this in turn is the condition for a more sound and fruitful development of individuals and society. An inadequate presentation of Christian anthropology gave rise to a wrong understanding of the relationship between human beings and the world. Often, what was handed on was a Promethean vision of mastery over the world, which gave the impression that the protection of nature was something that only the faint-hearted cared about. Instead, our “dominion” over the universe should be understood more properly in the sense of responsible stewardship.

Guidelines for this assignment

Follow these specifications:

1. For the purposes of this assignment, imagine that you are a newspaper or website news reporter who has been assigned to write a short article about the issues raised by Francis in your passage from Laudato Si’. Your job is not to “editorialize;” it is to report what Francis is trying to say, to the best of your understanding and to the best of your ability, but in your own words. (In other words, you cannot just quote whole chunks of your passage – you need to paraphrase and synthesize!) The readership of your paper or website consists of educated adults, but they are not necessarily experts on the details of the Rio Declaration (topic 1) and they might not know immediately what Francis means by “excessive anthropocentrism” or what it means to have a “Promethean vision of mastery over the world” (topic 2). So your job is in large part to explain, fill in details, and present the ideas Francis is discussing in a way that is more down-to-earth and less theoretical.

2. However, your presentation of Pope Francis’s statement should be specific, detailed, and correct.

3. If you use direct quotations from other sections of Laudato Si’, give the page and/or paragraph number(s) to identify the source. Here’s an example of what I’m looking for:

Francis clearly believes that companies should seek out sustainable uses of resources rather than pursuing only short-term profits, saying in §191 on page 140 of Laudato Si’ that, “efforts to promote a sustainable use of natural resources are not a waste of money, but rather an investment capable of providing other economic benefits in the medium term.”

4. If you consult any other sources – this means other books, web sites, etc. – and make use of direct quotations or ideas presented there but not in the passage, then provide enough information for an interested reader to find and check the source you are using. For books, give the author, title, publisher, year of publication, and page where your information is coming from. For web sites, give the full URL, an author if you can identify one, and the date you consulted the site.

5. The article (paper) as a whole should be well-organized and “flow” from one point to the next. You can start out with a few sentences about the general background of when this Encyclical Letter was written and what it is about in general terms. But do not spend more than two or three sentences on that. Do not repeat points unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

6. Your sentences should be clear, readable, and not overly wordy. You should aim for a lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

10. Errors in mechanics (spelling, punctuation, grammar) and usage should not annoy the reader (your editor). (A word of warning: Your editor tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid those mistakes.